One solution is 80% acid and another one is 30% acid. How much of each solution is needed to make a 150L solution that is 62% acid?
Found 2 solutions by fractalier, addingup:
Answer by fractalier(6550)

Call the amount of 80% acid used, x.
Then the amount of 30% acid used would be 150-x.
The setup looks like this:
(.80)x + (.30)(150-x) = (.62)(150)
Now solve this for x...
.8x + 45 - .3x = 93
.5x = 48
x = 96 L of 80% acid
Then 150 - x = 54 L of 30% acid

Answer by addingup(3677)

.8x+.3(150-x)= .62(150)
.8x+45-.3x= 93
.5x= 48
x= 96
You need 96 of the 80% solution and 150-96= 54 of the 30%
---------------------------------------------------------
Check - let's do a weighted average to prove this:
96/150= .64 (64% of the solution is at 80%)
54/150= .36 (36% of the solution is at 30%)
(.64*.8)+(.36*.3)= .62
.62= .62 We have the correct answer
J